# Sliced Rényi Pufferfish Privacy: Directional Additive Noise Mechanism and Private Learning with Gradient Clipping

### v2

@@ Appendix D Minimality and Attainability of HUC
Let $\mathsf{P}:=\gamma\times\mathbb{P}_{\eta,\rho}$ and $\mu_{t}:=(\Delta_{t})_{\#}\mathsf{P}$ be the pushforward measure of $\Delta_{t}$, and define
@@
where $\mathcal{B}_{\bar{\varepsilon}}(z)$ is a ball of radius $\bar{\varepsilon}$ centered at $z$. For each slice $u_{i}$, define
$\phi_{t}(u_{i}):=\sup_{\Delta\in\mathtt{ELS}_{t}}|\langle\Delta,u_{i}\rangle|\quad\text{and}\quad h^{\star}_{t,i}:=\phi_{t}(u_{i})^{2}.$
@@
    Attainability. If $\mathtt{ELS}_{t}$ is compact, then for each $i\in[m]$ there exists $\Delta_{i}^{\star}\in\mathtt{ELS}_{t}$ with $|\langle\Delta_{i}^{\star},u_{i}\rangle|=\phi_{t}(u_{i})$; hence $h^{\star}_{t,i}$ is attained.
However, the HUC constructed in Proposition 5.1 is generally not tight; it is a computable conservative bound. Proposition D.1 characterizes the minimal HUC and specifies when it is attained. In particular, any valid HUC $h_{t}$ must satisfy $h_{t,i}\geq h^{\star}_{t,i}$ for all $i\in[m]$, and the minimal cap $h_{t}^{\star}$ is attained whenever the essential shift set $\mathtt{ELS}_{t}$ is compact.
@@
The following corollary follows Proposition D.1.
Corollary D.1.
@@
D.1 Proof of Proposition D.1
We begin with a simple measure-theoretic observation. Let $\mu$ be a Borel probability measure on $\mathbb{R}^{d}$ with support $\mathsf{U}=\mathrm{supp}(\mu)$, and let $g:\mathbb{R}^{d}\to\mathbb{R}$ be continuous. If $g(\Delta)\leq c$ holds $\mu$-a.s., then necessarily
$\sup_{\Delta\in\mathsf{U}}g(\Delta)\leq c.$
Indeed, suppose for contradiction that there exists $\Delta^{\star}\in\mathsf{U}$ with $g(\Delta^{\star})>c$. By continuity of $g$, there exists $\bar{\varepsilon}>0$ such that $g(\Delta)>c$ for all $\Delta$ in the open Euclidean ball $B_{\bar{\varepsilon}}(\Delta^{\star})$. Since $\Delta^{\star}\in\mathsf{U}=\mathrm{supp}(\mu)$, we have $\mu\bigl(B_{\bar{\varepsilon}}(\Delta^{\star})\bigr)>0$, which contradicts the assumption $g(\Delta)\leq c$ $\mu$-a.s.
@@
In our setting, for any slice $u_{i}\in\mathbb{S}^{d-1}$ define the continuous map
$g_{u_{i}}(\Delta):=\bigl|\langle\Delta,u_{i}\rangle\bigr|,\qquad\Delta\in\mathbb{R}^{d}.$
The HUC inequality for a given $h_{t}$ asserts that
$g_{u_{i}}\bigl(\Delta_{t}(X,X^{\prime};R_{t})\bigr)=\bigl|\langle\Delta_{t}(X,X^{\prime};R_{t}),u_{i}\rangle\bigr|\;\leq\;\sqrt{h_{t,i}}\quad\text{$\mu_{t}$-a.s.},$
where $\mu_{t}$ is the pushforward measure of $\Delta_{t}$ and $\mathtt{ELS}_{t}=\mathrm{supp}(\mu_{t})$. Applying the auxiliary observation with $\mu=\mu_{t}$, $g=g_{u_{i}}$, and $c=\sqrt{h_{t,i}}$ yields
$\sup_{\Delta\in\mathtt{ELS}_{t}}\bigl|\langle\Delta,u_{i}\rangle\bigr|\;\leq\;\sqrt{h_{t,i}}.$
Hence $\phi_{t}(u_{i})\leq\sqrt{h_{t,i}}$ and therefore
$h^{\star}_{t,i}=\phi_{t}(u_{i})^{2}\leq h_{t,i}.$
This holds for each $i\in[m]$, proving minimality.
@@
Assume $\mathtt{ELS}_{t}$ is compact. The map $g_{u_{i}}(\Delta)=|\langle\Delta,u_{i}\rangle|$ is continuous, so by the Weierstrass Extreme Value Theorem it attains its maximum on $\mathtt{ELS}_{t}$. Thus there exists $\Delta^{\star}_{i}\in\mathtt{ELS}_{t}$ such that
$g_{u_{i}}(\Delta^{\star}_{i})=\max_{\Delta\in\mathtt{ELS}_{t}}g_{u_{i}}(\Delta)=\phi_{t}(u_{i}).$
Consequently, $h^{\star}_{t,i}=\phi_{t}(u_{i})^{2}$ is attained for each $i\in[m]$.
D.2 Proof of Corollary D.1
We work at a fixed iteration $t$ and suppress the $t$-index when the meaning is clear.
Step 1: Characterizing the feasible shifts. In the linear case,
@@
so for any coupled pair $(X,X^{\prime})$ and sampling draw $R_{t}$,
$\Delta_{t}(X,X^{\prime};R_{t})$ $=f_{t}(X,y_{<t};R_{t})-f_{t}(X^{\prime},y_{<t};R_{t})$
$=-A_{t}\,\bigl(\bar{g}_{t}(X;R_{t})-\bar{g}_{t}(X^{\prime};R_{t})\bigr).$
Let
$v:=\bar{g}_{t}(X;R_{t})-\bar{g}_{t}(X^{\prime};R_{t}).$
Under blockwise clipping with thresholds $\{C_{b}\}_{b=1}^{B_{\mathrm{blk}}}$ and discrepancy cap $K_{t}$ (i.e., at most $K_{t}$ indices differ in the minibatch), we have, for each block $b$,
$\bigl\|\mathsf{P}_{b}v\bigr\|_{2}\;\leq\;\frac{2C_{b}}{B_{t}}K_{t}\quad\text{$\gamma\times\mathbb{P}_{\eta,\rho}$-a.s.}$
since at most $K_{t}$ per-example contributions in that block can change, and each change is clipped to radius $C_{b}$ in that block. Thus any realized $v$ lies in the (blockwise) product of Euclidean balls
$\mathcal{V}:=\Bigl\{v\in\mathbb{R}^{d}:\bigl\|\mathsf{P}_{b}v\bigr\|_{2}\leq\alpha_{b}:=\frac{2C_{b}K_{t}}{B_{t}},\;b=1,\dots,B_{\mathrm{blk}}\Bigr\}.$
Consequently, any realized shift has the form
$\Delta=-A_{t}v\quad\text{with }v\in\mathcal{V}.$
By Proposition D.1, the minimal per-slice cap is
$h_{t,i}^{\star}=\sup_{\Delta\in\mathtt{ELS}_{t}}\bigl|\langle\Delta,u_{i}\rangle\bigr|^{2},$
where $\mathtt{ELS}_{t}$ is the essential shift set. Under the assumption that the blockwise gradient differences can attain the clipping radius in directions aligned with $\mathsf{P}_{b}A_{t}^{\top}u_{i}$, the essential shift set $\mathtt{ELS}_{t}$ coincides with the set
$\{-A_{t}v:v\in\mathcal{V}\},$
so
$h_{t,i}^{\star}=\sup_{v\in\mathcal{V}}\bigl|\langle-A_{t}v,u_{i}\rangle\bigr|^{2}=\sup_{v\in\mathcal{V}}\bigl|\langle v,A_{t}^{\top}u_{i}\rangle\bigr|^{2}.$
Step 2: Optimizing over the blockwise constraints. Fix a slice index $i$ and write
$w:=A_{t}^{\top}u_{i}.$
Decompose both $v$ and $w$ into blocks:
$v=\sum_{b=1}^{B_{\mathrm{blk}}}v_{b},\quad v_{b}:=\mathsf{P}_{b}v,\qquad w=\sum_{b=1}^{B_{\mathrm{blk}}}w_{b},\quad w_{b}:=\mathsf{P}_{b}w.$
The constraints on $v$ are
$\|v_{b}\|_{2}\leq\alpha_{b}:=\frac{2C_{b}K_{t}}{B_{t}},\qquad b=1,\dots,B_{\mathrm{blk}}.$
Then
$\langle v,w\rangle=\sum_{b=1}^{B_{\mathrm{blk}}}\langle v_{b},w_{b}\rangle.$
To maximize $\bigl|\langle v,w\rangle\bigr|$ subject to the blockwise constraints, we may choose each block $v_{b}$ independently. The optimal choice in block $b$ aligns $v_{b}$ with $w_{b}$ (or its negative), i.e.,
$v_{b}^{\star}=\alpha_{b}\frac{w_{b}}{\|w_{b}\|_{2}}\quad\text{whenever }w_{b}\neq 0,$
and $v_{b}^{\star}=0$ if $w_{b}=0$. This gives
$\langle v_{b}^{\star},w_{b}\rangle=\alpha_{b}\|w_{b}\|_{2},$
and the summed inner product becomes
$\sup_{v\in\mathcal{V}}\bigl|\langle v,w\rangle\bigr|=\sum_{b=1}^{B_{\mathrm{blk}}}\alpha_{b}\|w_{b}\|_{2}=\frac{2K_{t}}{B_{t}}\sum_{b=1}^{B_{\mathrm{blk}}}C_{b}\,\bigl\|\mathsf{P}_{b}A_{t}^{\top}u_{i}\bigr\|_{2}.$

### v3

@@ Appendix H Minimality and Attainability of HUC
For each direction $u_{i}$, define
$\phi_{t}(u_{i})\;:=\;\operatorname*{ess\,sup}_{(X,X^{\prime},R_{t})\sim\gamma\times\mathbb{P}_{\eta,\rho}}\big|\langle\Delta_{t}(X,X^{\prime};R_{t}),u_{i}\rangle\big|\textup{ and }h^{\star}_{t,i}:=\phi_{t}(u_{i})^{2}.$
In addition, let $\mathsf{P}:=\gamma\times\mathbb{P}_{\eta,\rho}$ and $\mu_{t}:=(\Delta_{t})_{\#}\mathsf{P}$ be the pushforward measure of $\Delta_{t}$, and define
@@
where $\mathcal{B}_{\bar{\varepsilon}}(z)$ is a ball of radius $\bar{\varepsilon}$ centered at $z$.
@@
    Attainability. If $\mathtt{ELS}_{t}$ is compact, then there exists $\Delta_{i}^{\star}\in\mathtt{ELS}_{t}$ such that $|\langle\Delta_{i}^{\star},u_{i}\rangle|=\sup_{\Delta\in\mathtt{ELS}_{t}}|\langle\Delta,u_{i}\rangle|$. Moreover, by definition of $\phi_{t}(u_{i})$ as an essential supremum, for very $\varepsilon>0$, $\mathsf{P}\!\left(\,|\langle\Delta_{t}(X,X^{\prime};R_{t}),u_{i}\rangle|\geq\phi_{t}(u_{i})-\varepsilon\,\right)>0$.
However, the HUC constructed in Proposition 5.1 is generally not tight; it is a computable conservative bound. Proposition H.1 characterizes the minimal HUC and specifies when it is attained. In particular, any valid HUC $h_{t}$ must satisfy $h_{t,i}\geq h^{\star}_{t,i}$ for all $i\in[m]$, and the minimal cap $h_{t}^{\star}$ is attained whenever the essential shift set $\mathtt{ELS}_{t}$ is compact.
@@
The following corollary follows Proposition H.1.
Corollary H.1.
@@
H.1 Proof of Proposition H.1
Let $\mathsf{P}:=\gamma\times\mathbb{P}_{\eta,\rho}$ and write
$Z_{i}(X,X^{\prime},R_{t})\;:=\;\big|\langle\Delta_{t}(X,X^{\prime};R_{t}),u_{i}\rangle\big|,\qquad i\in[m].$
Recall that $\phi_{t}(u_{i})=\operatorname*{ess\,sup}_{\mathsf{P}}Z_{i}$ and $h^{\star}_{t,i}=\phi_{t}(u_{i})^{2}$.
We use the following standard property: if $Z\geq 0$ is measurable and $Z\leq c$ holds $\mathsf{P}$-a.s., then $\operatorname*{ess\,sup}_{\mathsf{P}}Z\leq c$. Moreover, by definition of essential supremum, for every $\varepsilon>0$,
$\mathsf{P}\big(Z\geq\operatorname*{ess\,sup}_{\mathsf{P}}Z-\varepsilon\big)>0.$ (46)
@@
Let $h_{t}=(h_{t,1},\dots,h_{t,m})$ be any HUC for $(y_{<t},\gamma)$. By Definition 5.1, for each $i\in[m]$ we have
$Z_{i}(X,X^{\prime},R_{t})=\big|\langle\Delta_{t}(X,X^{\prime};R_{t}),u_{i}\rangle\big|\leq\sqrt{h_{t,i}}\quad\text{$\mathsf{P}$-a.s.}$
Applying the essential-supremum property gives
$\phi_{t}(u_{i})=\operatorname*{ess\,sup}_{\mathsf{P}}Z_{i}\leq\sqrt{h_{t,i}}.$
Squaring yields $h^{\star}_{t,i}=\phi_{t}(u_{i})^{2}\leq h_{t,i}$ for all $i\in[m]$, proving minimality.
@@
Define $\mu_{t}:=(\Delta_{t})_{\#}\mathsf{P}$ and $\mathtt{ELS}_{t}:=\mathrm{supp}(\mu_{t})$. Assume $\mathtt{ELS}_{t}$ is compact. The map $\Delta\mapsto|\langle\Delta,u_{i}\rangle|$ is continuous, hence by the Weierstrass Extreme Value Theorem there exists $\Delta_{i}^{\star}\in\mathtt{ELS}_{t}$ such that
$|\langle\Delta_{i}^{\star},u_{i}\rangle|=\sup_{\Delta\in\mathtt{ELS}_{t}}|\langle\Delta,u_{i}\rangle|.$
This establishes attainment of the set supremum over $\mathtt{ELS}_{t}$.
Finally, the essential attainability statement follows directly from (46) applied to $Z_{i}$: for every $\varepsilon>0$,
$\mathsf{P}\big(Z_{i}\geq\phi_{t}(u_{i})-\varepsilon\big)>0.$
Equivalently, in pushforward form,
$\mu_{t}\Big(\big\{\Delta:\ |\langle\Delta,u_{i}\rangle|\geq\phi_{t}(u_{i})-\varepsilon\big\}\Big)>0.$
H.2 Proof of Corollary H.1
Fix iteration $t$ and a slice $u_{i}\in\mathcal{U}$. Let $\mathsf{P}:=\gamma\times\mathbb{P}_{\eta,\rho}$. Write
@@
$\Delta_{t}(X,X^{\prime};R_{t})=-A_{t}\big(\bar{g}_{t}(X;R_{t})-\bar{g}_{t}(X^{\prime};R_{t})\big).$
Define
$v:=\bar{g}_{t}(X;R_{t})-\bar{g}_{t}(X^{\prime};R_{t}),\quad w:=A_{t}^{\top}u_{i}.$
Then $|\langle\Delta_{t},u_{i}\rangle|=|\langle v,w\rangle|$.
Step 1: a.s. feasible set for $v$.
Under blockwise clipping with thresholds $\{C_{b}\}_{b=1}^{B_{\mathrm{blk}}}$ and a discrepancy cap $K_{t}$, at most $K_{t}$ per-example contributions in the minibatch can differ between $X$ and $X^{\prime}$. Since each per-example block-gradient is clipped to $\ell_{2}$-norm at most $C_{b}$ in block $b$, the blockwise difference in the averaged clipped gradients satisfies, $\mathsf{P}$-a.s.,
$\|\mathsf{P}_{b}v\|_{2}\;\leq\;\frac{2C_{b}}{B_{t}}K_{t}\;=:\;\alpha_{b},\quad b=1,\dots,B_{\mathrm{blk}}.$
Equivalently, $v\in\mathcal{V}$ $\mathsf{P}$-a.s., where
$\mathcal{V}:=\Big\{v\in\mathbb{R}^{d}:\ \|\mathsf{P}_{b}v\|_{2}\leq\alpha_{b},\ \ b=1,\dots,B_{\mathrm{blk}}\Big\}.$
Step 2: upper bound on the minimal cap.
By Proposition H.1 (minimality with $\phi_{t}(u_{i})=\operatorname*{ess\,sup}_{\mathsf{P}}|\langle\Delta_{t},u_{i}\rangle|$),
$\phi_{t}(u_{i})=\operatorname*{ess\,sup}_{\mathsf{P}}|\langle v,w\rangle|\;\leq\;\sup_{v\in\mathcal{V}}|\langle v,w\rangle|.$
Now decompose $v=\sum_{b}v_{b}$ and $w=\sum_{b}w_{b}$ with $v_{b}:=\mathsf{P}_{b}v$, $w_{b}:=\mathsf{P}_{b}w$. Then, for any $v\in\mathcal{V}$,
$|\langle v,w\rangle|=\Big|\sum_{b=1}^{B_{\mathrm{blk}}}\langle v_{b},w_{b}\rangle\Big|\leq\sum_{b=1}^{B_{\mathrm{blk}}}\|v_{b}\|_{2}\,\|w_{b}\|_{2}\leq\sum_{b=1}^{B_{\mathrm{blk}}}\alpha_{b}\|w_{b}\|_{2}.$
Hence
$\sup_{v\in\mathcal{V}}|\langle v,w\rangle|=\sum_{b=1}^{B_{\mathrm{blk}}}\alpha_{b}\|w_{b}\|_{2}=\frac{2K_{t}}{B_{t}}\sum_{b=1}^{B_{\mathrm{blk}}}C_{b}\,\|\mathsf{P}_{b}A_{t}^{\top}u_{i}\|_{2}.$
@@
$\phi_{t}(u_{i})\;\leq\;\frac{2K_{t}}{B_{t}}\sum_{b=1}^{B_{\mathrm{blk}}}C_{b}\,\|\mathsf{P}_{b}A_{t}^{\top}u_{i}\|_{2}.$ (47)
Step 3: matching lower bound (attainability in essential-sup sense).
Consider the optimizer $v^{\star}\in\mathcal{V}$ defined blockwise by
$v_{b}^{\star}=\begin{cases}\alpha_{b}\,\dfrac{w_{b}}{\|w_{b}\|_{2}},&w_{b}\neq 0,\\[6.0pt] 0,&w_{b}=0,\end{cases}\quad b=1,\dots,B_{\mathrm{blk}}.$
Then $v^{\star}\in\mathcal{V}$ and
$\langle v^{\star},w\rangle=\sum_{b=1}^{B_{\mathrm{blk}}}\alpha_{b}\|w_{b}\|_{2}=\sup_{v\in\mathcal{V}}\langle v,w\rangle,$
so $\sup_{v\in\mathcal{V}}|\langle v,w\rangle|$ equals the right-hand side of (47).
By the corollary’s assumption (“the blockwise gradient differences can attain the clipping radius in directions aligned with $\mathsf{P}_{b}A_{t}^{\top}u_{i}$”), together with the requirement that the corresponding discrepancy pattern of size $K_{t}$ occurs with nonzero probability under the sampling rule, we have that for every $\varepsilon>0$,
$\mathsf{P}\Big(|\langle v,w\rangle|\geq|\langle v^{\star},w\rangle|-\varepsilon\Big)>0.$
By the defining property of essential supremum, this implies
$\phi_{t}(u_{i})=\operatorname*{ess\,sup}_{\mathsf{P}}|\langle v,w\rangle|=|\langle v^{\star},w\rangle|=\frac{2K_{t}}{B_{t}}\sum_{b=1}^{B_{\mathrm{blk}}}C_{b}\,\|\mathsf{P}_{b}A_{t}^{\top}u_{i}\|_{2}.$
Squaring yields